# Project page: https://xiaohanlei.github.io/projects/GaussNav/

### v2

@@ 3.3 Semantic Gaussian Construction
Differentiable Rendering via Splatting. Following the original 3DGS [15], our Semantic Gaussian renders an RGB image as follows: given a collection of 3D gaussians and camera pose, first sort all gaussians from front to back. RGB images can then be efficiently rendered by alpha-compositing the splatted 2D projection of each gaussian in order in pixel space. The rendered color of pixel ${\bf p}=(u,v)$ can be written as:
where $f_{i}(\bf{p})$ is computed as in Eq. (1) but with $\bm{\mu}$ and $r$ of the splatted 2D Gaussians in pixel-space:
@@
Following Keetha et al. [14], we also differentiably render depth and silhouette image, which determines the visibility. The depth image $D$ at pixel $\mathbf{p}$ is rendered as follows:
@@
Similarly, the silhouette image $S$ at pixel $\mathbf{p}$, which indicates the presence of Semantic Gaussian information, is rendered via:
@@
Gaussian Densification and Updating. Gaussian Densification aims to initialize new gaussians in the Semantic Gaussian at each new incoming frame. Following Keetha et al. [14], we add new gaussians based on a densification mask to determine which pixels should be densified:
$M(\textbf{p})=\Bigl{(}\hat{S}(\textbf{p})<0.5\Bigr{)}+\\ \Bigl{(}D(\textbf{p})<\hat{D}(\textbf{p})\Bigr{)}\Bigl{(}\textrm{L}_{1}\bigl{(}\hat{D}(\textbf{p})\bigr{)}>50\textrm{MDE}\Bigr{)},$ (7)
where the first term represents where the Semantic Gaussian isn’t adequately dense, and the second term indicates where the ground-truth depth is in front of the predicted depth and the depth error is greater than 50 times the median depth error (MDE). After densifying current Semantic Gaussian, we update the parameters of gaussians given poses and observations. This is done by differentiable-rendering and gradient-based-optimization, which is equivalent to the “classic” problem of fitting a radiance field to images with known poses.
3.4 Gaussian Navigation
Once the Semantic Gaussian is constructed and the agent’s initial pose is known, transforming IIN into PointGoal Navigation requires: 1. converting the Semantic Gaussian into a map form that can be used for path planning, i.e., either 3D voxels $M_{3D}$ or 2D BEV grids $M_{2D}$, and 2. predicting the position of the target object instance $\hat{P}_{g}$. Therefore, we first convert the Semantic Gaussian into a point cloud, whereby each gaussian is reduced to a single point in the point cloud representation. The point cloud is then voxelized into 3D voxels $M_{3D}$, then the 3D voxels $M_{3D}$ are projected to 2D BEV grids $M_{2D}$. By classifying and matching in the pre-built Semantic Gaussian $G_{s}$ with goal image $I_{g}$, we can ground target object’s position $\hat{P}_{g}$, as depicted in Fig. 2(b). Given start pose $P_{0}$, predicted target object’s position $\hat{P}_{g}$ and grid map $M_{2D}$, we can easily use Fast Marching Method (FMM)  [^1] https://pythonhosted.org/scikit-fmm/ to compute the shortest path and the corresponding action $a_{t}$.
@@ 3.4 Gaussian Navigation
Match & Grounding. With the predicted target object’s label $\hat{l}_{g}$, we identify all candidate objects that share the same class label. For each candidate instance, we generate a set of descriptive images by rendering the object from multiple viewpoints to capture its features. $n$ denotes the observed instances of the same class as target object, $S_{i}$ indicates the $i$-th object instance’s descriptive renderings, then the universal set of $S$ can be formulated as:
@@
$i_{\text{max}}=arg\max\{\max_{s\in S_{1}}K(s),\max_{s\in S_{2}}K(s),\ldots,\max_{s\in S_{n}}K(s)\}\quad\text{for}\quad i=1,2,\ldots,n,$ (9)
@@
where $K(\cdot)$ is defined as the matched number of keypoints between renderings and goal image $I_{g}$. Specifically, we extract the pixel-wise $(x,y)$ coordinates of the keypoints and their associated feature descriptors using DISK [33] both on RGB renderings and goal image. Subsequently, the matched pairs are computed using LightGlue [23]. The candidate object whose rendered images yield the highest number of matched keypoints is then selected as the target object, as in Eq. (9).
@@
When the object instance is selected, we ground the object both in the 3D voxels $M_{3D}$ and 2D BEV grid map $M_{2D}$. Due to the presence of outliers in the projection of an object instance onto the 2D BEV map, which are caused by errors in semantic segmentation, we perform clustering on the instances on the map. Specifically, we use Density-based Spatial Clustering of Applications with Noise (DBSCAN)  [^2] https://scikit-learn.org/stable/modules/generated/sklearn.cluster.DBSCAN.html, which groups together points that are closely packed together, marking as outliers points that lie alone in low-density regions. With the precise selected object instance’s location, we can easily transform the IIN task to a PointGoal task.
@@
Path Planning. Given 2D BEV grid map $M_{2D}$, along with the agent’s starting position and the goal’s location, we can efficiently generate a shortest distance field using FMM. Each point within this field encapsulates the minimal distance necessary to traverse from the starting point to the goal. We extract a relevant subset of this distance field that falls within the agent’s operational range. Subsequently, a waypoint is chosen from this subset to ensure it avoids any intersections with obstacles while adhering to a local minimum in the distance field. With the selected waypoint, the agent can readily calculate an action based on the angle and distance from its current state. The agent iterates the aforementioned process to generate a sequence of actions, continuing until the destination is reached.
4 Experiments
4.1 Experiment Setup
Evaluation Metrics. We evaluate our model with both success and efficiency. We report Success Rate (Success), Success rate weighted by normalized inverse Path Length (SPL). An episode is deemed successful ($\text{Success}=1$) if the agent invokes the STOP action within a $1.0$m Euclidean distance from the goal object. SPL is an efficiency measure defined in [2], is given by:
$\text{SPL}=\frac{1}{N}\sum_{i=1}^{N}S_{i}\cdot\frac{l_{i}}{\max(p_{i},l_{i})},$ (10)
4.2 Comparison to the state-of-the-art methods
We evaluate our proposed model against various baselines and previous state-of-the-art work, as presented in Tab. 1.
Table 1: Performance of our proposed GaussNav as compared to the baselines and previous state-of-the-art methods on the HM3D [41] datasets across two different metrics: Success and SPL [2].
@@ 4.2 Comparison to the state-of-the-art methods
RL Baseline. The baseline model for the IIN task is trained end-to-end using Proximal Policy Optimization (PPO). Training is conducted from scratch as described by Schulman et al. [31]. Despite requiring a substantial 3.5 billion steps for the RL agent to converge, the ability to generalize to unseen scenes is lacking (row 1 in Tab. 1).
@@
OVRL-v2 Baseline. Offline Visual Representation Learning v2 (OVRL-v2) was introduced by Karmesh et al. [39]. This novel approach to model-free, end-to-end navigation policies underscores the importance of self-supervised pretraining for visual encoders and presents the innovation of a Compression Layer (CL). Originally, OVRL-v2 was trained for the ImageGoal Navigation task. Direct application of OVRL-v2 to IIN task without fine-tuning yields suboptimal performance, as indicated by a Success of 0.006 (row 2 in Tab. 1). This reduced efficacy can be attributed to several factors: the transition between scene datasets from Gibson [37] to HM3D [41], differences in robot embodiment from Locobot  [^4] http://www.locobot.org/ to Stretch, and a shift in the nature of goal destinations from image sources to image subjects. Fine-tuning OVRL-v2 specifically for IIN task on the HM3D dataset significantly improves outcomes, resulting in a Success of 0.248 (row 3 in Tab. 1).
Mod-IIN. Mod-IIN [17] presents a decomposition of the IIN task into four sub-tasks: exploration, goal instance re-identification, goal localization, and local navigation. This method utilizes feature matching to re-identify the goal instance within the egocentric vision and projects the matched features onto a map to localize the goal. Each sub-task is addressed using off-the-shelf components that do not require any fine-tuning. It achieves a Success of 0.561 and SPL of 0.233 (row 4 in Tab. 1).
IEVE. The IEVE approach [21] is inspired by the human heuristic of “getting closer to confirm” when attempting to distinguish objects from a distance. IEVE adopts a modular architecture that enables dynamic switching between exploration, verification, and exploitation actions. This flexibility empowers the agent to make informed decisions tailored to varying circumstances. Results demonstrate that IEVE, when integrated with Mask RCNN [12], achieves Success of 0.684 and an SPL of 0.241, as well as Success of 0.702 and SPL of 0.252 (rows 5 and 6 in Tab. 1) when using InternImage [34].
GaussNav (ours). GaussNav demonstrates a significant performance advantage over all baseline models (row 7 in Tab. 1). It significantly surpasses all existing models in terms of SPL by a huge margin of 0.326 (rows 6 and 7 in Tab. 1). We attribute the superior performance of our method to the direct mapping from goal image to goal position, thereby facilitating efficient navigation without additional exploration. In its initial episode within a new scene, agent explores and construct the Semantic Gaussian of the scene. Semantic Gaussian enables agent to efficiently and accurately locate the goal’s location in following episodes.
4.3 Ablation Study
Table 2: Ablations. Table showing comparison of our proposed model, GaussNav, with ablations and ground truth goal position and Match on HM3D dataset.
@@ 4.3 Ablation Study
GaussNav w.o. Classifier. In Fig. 2(b), we replace the Classifier’s output with a random generated target category. We observe that the Success drops to 0.375 and the SPL decreases to 0.291 (row 2 in Tab. 2). This is because correct classification is a prerequisite for our method to filter potential candidates.
GaussNav w.o. Match. The Match module is designed to distinguish target from candidates with the same class. Without the Match module, we randomly select from these candidates. The Success and SPL falls to 0.444 and 0.353 (row 3 in Tab. 2). This can be attributed to the random selection of candidate instances without Match module.
GaussNav w. SIFT or GlueStick [28]. To evaluate the impact of different extractors and matching algorithms on our GaussNav, we replace the combination of DISK [33] + LightGlue [23]. We employ both SIFT + FLANN and GlueStick [28] as alternatives. The first replacement represents a greater decrease, (row 4 in Tab. 2) and the second only displays a slight decrease (row 5 in Tab. 2). These results demonstrate the varying performance of different local feature matching algorithms on the HM3D dataset.
4.4 Efficiency Analysis
Table 3: Number of object instances. The table displays the number of object instances across different categories in the first floor of scene CrMo8WxCyVb.
@@ 4.4 Efficiency Analysis
Locating the target object within the map is a critical component of our method, which leverages renderings to match and ground the goal object effectively. A naive approach that entails comparing all possible observations at each navigable viewpoint is computationally prohibitive. To address this challenge, we introduce an optimization that constructs semantic gaussians, allowing us to group object instances under their corresponding semantic labels.
Our method significantly reduces the overhead by limiting the comparisons to several descriptive renderings of each instance, rather than searching the entire space. For instance, consider the scene CrMo8WxCyVb-floor--2.9. The navigable area on the first floor of this scene is quantified to be $54.03m^{2}$. This space can be discretized into approximately 54 squares, each with an area of $1m^{2}$. Positioned within each square, the agent can observe its surroundings from 12 distinct viewing angles, covering a full $360\deg$ with each angle spanning $30\deg$. Thus, the original search space for this single floor would consist of $12\times 54=648$ potential observations.

### v3

@@ III-C Semantic Gaussian Construction
The $\bm{\mu}$ and $r$ are the splatted 2D Gaussians in pixel-space:
@@
Render. Different from 3DGS [29], we differentiably render depth and silhouette image, which determines the visibility and will be used for the next Gaussian Densification and Updating. The depth $D$ and silhouette image $S$ at pixel $\mathbf{p}$ is rendered as follows:
@@
Semantic Segmentation. As depicted in Figure 4, the agent’s RGB observation $I_{t}$ is segmented into $C_{t}$ using Mask-RCNN [30]. The segmented result will be used for initializing new Gaussians in Gaussian Densification and supervising the Gaussians’ parameters in Semantic Gaussian Updating.
Gaussian Densification. Gaussian Densification is performed by comparing the rendered results at position $P_{t}$ using Gaussians from $t-1$ with the ground truth. This process adds new Gaussians where the previous Gaussians fail to represent the scene in the new observations. Following Keetha et al.[43], we add new Gaussians based on a densification mask to determine which pixels should be densified:
$\small M(\textbf{p})=\Bigl{(}\hat{S}(\textbf{p})<0.5\Bigr{)}+\Bigl{(}D(\textbf{p})<\hat{D}(\textbf{p})\Bigr{)}\Bigl{(}\textrm{L}_{1}\bigl{(}\hat{D}(\textbf{p})\bigr{)}>50\textrm{MDE}\Bigr{)},$ (7)
where the first term represents where the Semantic Gaussian is not adequately dense, and the second term indicates where the ground-truth depth is in front of the predicted depth and the depth error is greater than 50 times the median depth error (MDE).
Semantic Gaussian Updating. After densifying current Semantic Gaussian, we update the parameters of Gaussians given poses and observations. This is done by differentiable-rendering and gradient-based-optimization, which is equivalent to the “classic” problem of fitting a radiance field to images with known poses. Specifically, we update the parameters of Gaussians by minimizing the RGB, depth and segmentation errors.
Figure 4: An illustration of Semantic Gaussian Construction. At timestep $t$, the pipeline updates the Gaussians from $t-1$ through densification and updating, which involves a comparison between the rendered RGB and depth images against the current input training views. Concurrently, semantic labels are assigned to the densified Gaussians using the segmented images. Finally, the Gaussians are refined through differentiable rendering.
Figure 5: An illustration of Gaussian Navigation. Our approach begins with the classification of the goal image using pre-constructed Semantic Gaussian. Upon determining the predicted class, we generate descriptive images around instances belonging to that class. These images are then matched with the target object to identify and ground the goal instance. Utilizing the map and the established goal, the agent employs path planning to compute the sequence of actions.
III-D Gaussian Navigation
To navigate using a constructed Semantic Gaussian, we propose Gaussian Navigation, as illustrated in Figure 5. We first classify the goal image $I_{g}$ to predict the semantic label $\hat{l}_{g}$ of the current target, such as ‘chair’ in Figure 5. This semantic label is then used to query the relevant Gaussians. For each relevant object instance of the same class, we render descriptive images. These renderings are compared with the goal image to ground the goal object, predicting its position $\hat{P}_{g}$. Finally, the Path Planning module generates a feasible path and determines the agent’s action.
@@ III-D Gaussian Navigation
Match & Grounding. With the predicted target object’s label $\hat{l}_{g}$, we identify all candidate objects that share the same class label. For each candidate instance, we generate a set of descriptive images by rendering the object from multiple viewpoints to capture its features. Specifically, for one training view containing possible candidate objects, we augment it to $n_{v}$ views by novel view synthesis (NVS). we denote the transformation from the camera to the world coordinate system of the training view as c2w, and record the translation of the potential target object in the camera frame as $\mathbf{t}^{o}_{c}$. We define the rotation matrix from the object to the world frame using the forward, right, and up vectors. The forward direction is the vector from the training view to the potential target object, namely $\mathbf{t}^{o}_{c}$. The up direction is the vector [0,1, 0]01 0[0,\,-1,\,0][ 0 , - 1 , 0 ], and the right vector is orthogonal to both the forward and up vectors, forming a right-handed coordinate system with the right, up, and forward vectors. The translation vector from the object to the world frame is
$\mathbf{t}^{o}_{w}=\text{c2w}\times\mathbf{t}^{o}_{c}.$ (8)
Together, the rotation matrix and translation vector constitute the rigid transformation matrix o2w from the object to the world frame.
We define the rotation matrices around the $y$-axis and $x$-axis as $\mathbf{R}_{y}(\theta)$ and $\mathbf{R}_{x}(\theta)$, respectively, representing new viewpoints formed by rotating around the object by an angle $\theta$ in the horizontal and vertical directions. Thus, the transformation from the camera to the world frame for new viewpoints in the horizontal direction is
[…]
and in the vertical direction:
[…]
In experiments, when $n_{v}=1$, we do not perform NVS; when $n_{v}=3$, we perform NVS at $\theta=\pm 15^{\circ}$ (both horizontal and vertical); and when $n_{v}=5$, we use $\theta=\pm 15^{\circ}$, ±$\pm 30^{\circ}$.
After NVS, the original training views are augmented. The augmented rendering set for the $i$-th object instance is denoted as $S_{i}$. Let $n$ denote the observed instances of the same class as target object, then the universal set of $S$ can be formulated as:
@@
$\begin{split}i_{\text{max}}=arg\max\{\max_{s\in S_{1}}\Omega(s),\max_{s\in S_{2}}\Omega(s),\ldots,\max_{s\in S_{n}}\Omega(s)\}\\ \text{for}\quad i=1,2,\ldots,n,\end{split}$ (12)
@@
where $\Omega(\cdot)$ is defined as the matched number of keypoints between renderings and goal image $I_{g}$. Specifically, for the rendering $s\in S_{i}$ of the $i$-th object instance and the goal image $I_{g}$, we extract the pixel-wise $(x,y)$ coordinates of the keypoints and their associated feature descriptors $V_{t}$ using DISK [78]. That is:
$(K_{t},V_{t})=\text{DISK}(s),\quad(K_{g},V_{g})=\text{DISK}(I_{g}).$ (13)
Subsequently, the matched pairs $(\hat{K_{t}},\hat{K_{g}})$ are computed using LightGlue [79]. The feature matching process is formulated as follows:
$(\hat{K_{t}},\hat{K_{g}})=\text{LightGlue}((K_{t},V_{t}),(K_{g},V_{g})).$ (14)
Thus, the number of matched points, i.e., the length of $\hat{K_{t}}$ or $\hat{K_{g}}$, is denoted as $\Omega$. The candidate object whose rendered images yield the highest number of matched keypoints is then selected as the target object, as shown in Equation 12.
@@
When the object instance is selected, we ground the object in the Semantic Gaussian. Due to the presence of outliers, which are caused by errors in semantic segmentation, we perform clustering on the instances on the map. Specifically, we use Density-based Spatial Clustering of Applications with Noise (DBSCAN)  [^1] https://scikit-learn.org/stable/modules/generated/sklearn.cluster.DBSCAN.html, which groups together points that are closely packed together, marking as outliers points that lie alone in low-density regions. With the precise selected object instance’s location, we can easily transform the IIN task to a PointGoal task.
Path Planning. The Semantic Gaussian is not designed for path planning. We first convert the Semantic Gaussian into a point cloud, whereby each Gaussian is reduced to a single point in the point cloud representation. The point cloud is then voxelized into 3D voxels $M_{3D}$, then the 3D voxels $M_{3D}$ are projected to 2D BEV grids $M_{2D}$. Here, we use the 2D projection of Semantic Gaussian rather than the 2D geometric map used in pre-exploration to maintain a consistent representation throughout the pipeline. Our Semantic Gaussian initializes Gaussians using point cloud derived from depth image and does not prune any Gaussian in the optimization stage. Thus, 2D projection of Semantic Gaussian or 2D geometric map are fundamentally equivalent.
@@
Given 2D BEV grid map $M_{2D}$, along with the agent’s starting position and the goal’s location, we can efficiently generate a shortest distance field using FMM. Each point within this field encapsulates the minimal distance necessary to traverse from the starting point to the goal. We extract a relevant subset of this distance field that falls within the agent’s operational range. Subsequently, a waypoint is chosen from this subset to ensure it avoids any intersections with obstacles while adhering to a local minimum in the distance field. With the selected waypoint, the agent can readily calculate an action based on the angle and distance from its current state. The agent iterates the aforementioned process to generate a sequence of actions, continuing until the destination is reached.
IV Experiments
This section first details our experiment setup, followed by a comparative analysis with state-of-the-art approaches. Subsequently, we present an ablation study to evaluate the efficacy of individual components in our proposed methods. Finally, we assess the computational efficiency of our approach, examine error cases, and evaluate the rendering quality.
@@ IV Experiments
FGPrompt [80] 0.099 0.028
MultiON Baseline [51] 0.066 0.045
MultiON Implicit [53] 0.143 0.107
MultiON Camera [50] 0.186 0.142
@@
Mod-IIN [20] (Scene Map) 0.563 0.323
IEVE Mask RCNN [22] (Scene Map) 0.683 0.331
IEVE InternImage [22] (Scene Map) 0.705 0.347
@@
TABLE I: Performance comparison of our GaussNav with previous state-of-the-art methods on the HM3D [14] datasets across two different metrics: Success and SPL [81]. The table is divided into four sections. The first section presents the results of end-to-end methods. The second section shows the transfer performance of MultiON-related methods on the IIN task. The third section includes the state-of-the-art methods on the IIN task. Finally, the fourth section aims to provide a fair comparison with GaussNav by replacing the episodic map used in these methods from the third section with a scene-specific map, allowing the agent to retain the map from the previous episode.
IV-A Experiment Setup
Figure 6: Trajectory Analysis. Our Semantic Gaussian map representation can allow agent to directly ground target object from a single goal image, thereby facilitating efficient visual navigation.
Evaluation Metrics. Following Krantz et al.[19], we evaluate our model with both success and efficiency. We report Success Rate (Success), Success rate weighted by normalized inverse Path Length (SPL). An episode is deemed successful ($\text{Success}=1$) if the agent invokes the STOP action within a $1.0$m Euclidean distance from the goal object. SPL is an efficiency measure defined in [81], is given by:
$\text{SPL}=\frac{1}{N}\sum_{i=1}^{N}S_{i}\cdot\frac{l_{i}}{\max(p_{i},l_{i})},$ (15)
IV-B Comparison to the State-of-the-art Methods
We evaluate our proposed model against various baselines and previous state-of-the-art work, as presented in Table I. Unlike conventional IIN methods that generate a map for each episode, GaussNav constructs a map across episodes in a scene. The first block lists results from end-to-end baselines, and the second block includes methods designed for the MultiON task but implemented here for the IIN task. The third block of Table I presents the performance of original state-of-the-art IIN methods, while the fourth block reports their performance when adapted to use scene map.
End-to-end Baselines. We evaluate the performance of two end-to-end methods on the IIN task. RL Baseline built a network that observes agent RGB $(\mathcal{V}_{RGB})$, agent depth $(\mathcal{V}_{D})$, the goal image $(\mathcal{V}_{G})$, GPS coordinates $(x,z)$, and heading $(\theta)$. Visual observations are encoded with separate ResNet-18 [82] encoders and GPS and heading are encoded with 32-dimensional linear layers. Then, these features are concatenated and encoded with a 2-layer LSTM. Finally, an action $a(t)$ is sampled from a categorical distribution. The network was trained from scratch using Proximal Policy Optimization (PPO) [83].
@@ IV-B Comparison to the State-of-the-art Methods
OVRL-v2, proposed by Yadav et al.[55], introduced self-supervised pretraining for visual encoders in ImageGoal Navigation. Originally, OVRL-v2 was trained for the ImageGoal Navigation task. Direct application of OVRL-v2 to IIN task without fine-tuning yields suboptimal performance, as indicated by a Success of 0.006 (row 2 in Table I). This reduced efficacy can be attributed to several factors: the transition between scene datasets from Gibson [13] to HM3D [14], differences in robot embodiment from Locobot  [^3] http://www.locobot.org/ to Stretch, and a shift in the nature of goal destinations from image sources to image subjects. Fine-tuning OVRL-v2 specifically for IIN task on the HM3D dataset significantly improves outcomes, resulting in a Success of 0.248 (row 3 in Table I).
State-of-the-art Methods in MultiON. The MultiON task shares many similarities with the Scene-specific Map representation we employ in evaluating GaussNav. We re-implement several state-of-the-art methods [51, 50, 53] from the MultiON task on the IIN benchmark. It is worth noting that we directly input the semantic category of the target object to the aforementioned methods. As the navigation requirements of the IIN task are at the instance level rather than the semantic level, this discrepancy in task formulation leaves room for performance improvement.
@@
TABLE II: Ablation study of GaussNav. We study the impact of Classifier, Match module, novel view synthesis (NVS) and different local feature extraction and matching algorithms on our GaussNav. The last two rows describe the performance of our method using ground truth match results and goal position.
Method Metric Ch. So. Pl. Bed Tol. TV
w.o. clf Success $\uparrow$ 0.821 0.873 0.798 0.914 0.936 0.829
Time (s) $\downarrow$ 30.9 24.8 33.2 18.1 24.6 31.9
w. clf Success $\uparrow$ 0.782 0.859 0.854 0.878 0.702 0.847
Time (s) $\downarrow$ 15.7 8.09 11.6 5.97 1.85 2.74
TABLE III: Results of matching success and time taken with or without classifier.(Abbreviations: clf = classifier, Ch. = Chair, So. = Sofa, Pl. = Plant, Tol. = Toilet)
State-of-the-art Methods in IIN. For fair comparison, We evaluate the performance of previous state-of-the-art methods [20, 22] on the IIN task using two types of map representations: episodic map and scene-specific map. Mod-IIN [20] decomposes the IIN task into exploration, goal instance re-identification, goal localization, and local navigation. This method utilizes feature matching to re-identify the goal instance within the egocentric vision and projects the matched features onto a map to localize the goal. Each sub-task is addressed using off-the-shelf components that do not require any fine-tuning.
IEVE [22] employs a modular architecture that dynamically switches between exploration, verification, and exploitation actions. This flexibility empowers the agent to make informed decisions tailored to varying circumstances. Mod-IIN and IEVE both demonstrate exceptional performance on the IIN task. We also implement the scene-specific map representation in these methods for a fair comparison with GaussNav. As shown in Table I, GaussNav demonstrates a significant performance advantage over all methods. It significantly surpasses all existing models in terms of SPL by a huge margin of 0.231 (last 2 rows in Table I). This result indicates that our Semantic Gaussian map, by preserving the intricate texture details of objects in the scene, enables the agent to directly locate the target object based on the goal image without the need for additional verification.
We attribute the superior performance to the novel map representation of Semantic Gaussian, allowing agent to directly ground goal target without additional exploration or verification, as evidenced by Figure 6. Unlike previous widely-used BEV map representation, our Semantic Gaussian can allow agent to query Gaussian through semantic label input and render descriptive images of an object instance. Therefore, our GaussNav does not require explicit verification of potential object candidates, unlike BEV map-based approaches. Instead, it selects the most probable candidate from a multitude of possibilities and navigates towards it directly.
Figure 7: SPL and FPS Analysis. Red markers represent end-to-end methods, while blue markers indicate modular approaches. Our GaussNav belongs to the modular approach, achieving the highest frame rate among modular methods while attaining the highest SPL across all approaches in IIN task.
Figure 8: Visualization of novel view synthesis results. The visualization results include horizontal and vertical novel view synthesis results with $\theta=\pm 15^{\circ}$. The upper right corner of each image shows the number of matched keypoints with the goal image.
@@
TABLE IV: Number of object instances across different categories in the first floor of scene CrMo8WxCyVb.
Metric original GT original horizontal vertical GT horizontal GT vertical
$n_{v}=1$ $n_{v}=1$ $n_{v}=3$ $n_{v}=5$ $n_{v}=3$ $n_{v}=5$ $n_{v}=3$ $n_{v}=5$ $n_{v}=3$ $n_{v}=5$
Success $\uparrow$ 0.811 0.815 0.824 0.827 0.832 0.831 0.835 0.846 0.839 0.845
Time (s) $\downarrow$ 11.5 11.7 32.1 53.9 33.1 54.2 32.7 54.1 33.4 55.1
TABLE V: Results of the number of rendered images $n$, different directions (vertical or horizontal) and whether to use ground truth rendering’s impact on the matching success.
IV-C Ablation Study
GaussNav w.o. Classifier. In Figure 5, we replace the Classifier’s output with a random generated target category. We observe that the Success drops to 0.375 and the SPL decreases to 0.291 (row 2 in Table II). To better evaluate the classifier, we design the following experiment. We define the training views from the full space of candidate objects as the complete set, and the training views filtered by the classifier as a subset. Given the goal image and the keypoint matcher, we compute, in each set respectively, the view with the largest number of matching keypoints. If this view contains the goal object, we consider it a success; otherwise, a failure. We also record the total time spent in the entire local feature matching process to assess the impact of the classifier on efficiency. Here, we do not use the navigation metrics such as Success and SPL because this allows us to eliminate the influence of irrelevant factors, such as path planning errors. Our experimental results are shown in Table III. As can be seen, the total success without the classifier is slightly higher than that with the classifier (an increase of 0.039), but the improvement is limited. However, the time taken is 2.5 times that with the classifier, making it significantly less efficient. Therefore, considering the trade-off between performance and efficiency, we choose to include the classifier as a component of GaussNav framework.
GaussNav w.o. Match. The Match module is designed to distinguish the target from candidates with the same class. Without the Match module, we randomly select from these candidates. The Success and SPL falls to 0.444 and 0.353 (row 3 in Table II). This can be attributed to the random selection of candidate instances without Match module.
GaussNav w. SIFT or GlueStick [84]. To evaluate the impact of different extractors and matching algorithms on our GaussNav, we replace the combination of DISK [78] + LightGlue [79]. We employ both SIFT + FLANN and GlueStick [84] as alternatives. The first replacement represents a greater decrease, (row 4 in Table II) and the second only displays a slight decrease (row 5 in Table II). These results demonstrate the varying performance of different local feature matching algorithms on the HM3D dataset.
NVS Analysis. We conduct experiments to evaluate the impact of the number of rendered images $n_{v}$ and the upper bound of NVS. Specifically, we perform ablation studies on NVS with $n_{v}=3$ and $n_{v}=5$ in both horizontal and vertical directions, and also evaluate the upper bound achievable using GT rendered results. To avoid the influence of other factors in navigation, we consider it a success only if the image retrieved through keypoint matching contains the target object; otherwise, it is considered a failure. The specific evaluation metrics are detailed in GaussNav w.o. Classifier.. The experimental results are presented in Table V. Overall, utilizing NVS is beneficial for successfully recognizing objects. However, a large $n_{v}$ does not necessarily yield positive effects (see Table V, vertical $n_{v}=5$ vs. vertical $n_{v}=3$). This is because our Semantic Gaussian map may have only a few observational viewpoints for a particular object, unlike the dozens available in traditional 3DGS. Therefore, when rendering from novel viewpoints, artifacts such as holes may occur, reducing the number of features, as visualized in Figure 8. Nevertheless, for GT NVS, the improvement is significant.
Method Ver. NVS Hor. NVS GT Success $\uparrow$ SPL $\uparrow$
IEVE [22] - - - 0.702 0.252
GaussNav (ours) $\checkmark$ $\times$ $\times$ 0.713 0.259
GaussNav (ours) $\times$ $\checkmark$ $\times$ 0.715 0.261
GaussNav (ours) $\checkmark$ $\checkmark$ $\times$ 0.723 0.265
GaussNav (ours) $\checkmark$ $\checkmark$ $\checkmark$ 0.747 0.289
TABLE VI: Performance Comparison of IEVE [22] and GaussNav. We compare GaussNav with the state-of-the-art method IEVE using different map representations. (Abbreviations: Ver. = Vertical, Hor. = Horizontal)
To further study the impact of NVS without pre-exploration, we adopt the same framework as IEVE [22], with the only difference being the map representation. IEVE uses 2D BEV map while GaussNav implements Semantic Gaussian map. Additionally, GaussNav performs NVS whenever a similar object appears in each observation. The results are presented in Table VI. Our Semantic Gaussian map enhances the existing observations without relying on pre-exploration. While the improvement may not be as prominent as the ability of instance-level object localization, the result of last row in Table VI indicates promising direction for future work.
IV-D Efficiency Analysis
Figure 9: Qualitative example of our GaussNav agent performing IIN task in the Habitat simulator. When randomly initialized in the environment, the agent is given a goal image depicted a target object. Gaussian Navigation directly grounds the target object in the Semantic Gaussian and guide the agent towards it.
Figure 10: Semantic Gaussian visualization results.
We analyze the temporal efficiency of our approach from two perspectives: preprocessing and agent-environment interaction. In the context of our method, “preprocessing” refers to the process of matching and grounding target objects given the Semantic Gaussian. The latter perspective compares the runtime frame rate of our method with various other approaches.